Integrating Factors found by Inspection.


1. y(2xy + 1)dx − xdy = 0


2. y(x^4 − y^2)dx + x (x^4 + y^2)dy = 0


3. (x^3y^3 + 1)dx + x^4y^2dy = 0


4. y(x^2y^2 − 1)dx + x(x^2y^2 + 1)dy = 0


5. y(2x + y^2)dx + x(y^2 − x)dy = 0


6. y(3x^3 − x + y)dx + x^2(1 − x^2)dy = 0


7. y(3x^3 − x + y)dx + x^2(1 − x^2)dy = 0


8. y^2(1 − x^2)dx + x(x^2y + 2x + y) = 0


1
Expert's answer
2021-03-29T17:03:54-0400

Solution.

1. "y(2xy + 1)dx \u2212 xdy = 0"

"2xdx+\frac{ydx-xdy}{y^2}=0"

"2xdx+d(\frac{x}{y})=0"

"\int 2xdx+\int d(\frac{x}{y})=0"

"x^2y+x=Cy," where "C" is some constant.

2. "y(x^4 \u2212 y^2)dx + x (x^4 + y^2)dy = 0"

"x^4(ydx+xdy)+y^2(xdy-ydx)=0"

"ydx+xdy+\frac{y^2}{x^2}\frac{xdy-ydx}{x^2}=0"

"d(xy)+(\frac{y}{x})^2d(\frac{y}{x})=0"

"\int d(xy)+\int(\frac{y}{x})^2d(\frac{y}{x})=0"

"3x^4y+y^3=Cx^3," where "C" is some constant.

3. "(x^3y^3 + 1)dx + x^4y^2dy = 0"

"x^3y^2(ydx+xdy)+dx=0"

"x^2y^2d(xy)+\frac{dx}{x}=0"

"x^3y^3+3\ln{x}=C," where "C" is some constant.

4. "y(x^2y^2 \u2212 1)dx + x(x^2y^2+1)dy = 0"

"x^2y^3dx-ydx+x^3y^2dy+xdy=0"

"\frac{xdy-ydx}{x^2}+y^2(ydx+xdy)=0"

"d(\frac{y}{x})+y^2d(xy)=0
\newline
\frac{x}{y}d(\frac{y}{x})+xyd(xy)=0"

"2\ln{\frac{y}{x}}+x^2y^2=C," where "C" is some constant.

5. "y(2x + y^2)dx + x(y^2 \u2212 x)dy = 0"

"2xydx+y^3dx+xy^2dy-x^2dy=0"

"\frac{2xydx-x^2dy}{y^2}+ydx+xdy=0"

"d(\frac{x^2}{y})+d(xy)=0"

"x^2+xy^2=Cy," where "C" is some constant.

6. and 7.

"y(3x^3 \u2212 x + y)dx + x^2(1 \u2212 x^2)dy = 0"

"x^3-x=Cy-y\ln{x}," where "C" is some constant.

8. "y^2(1 \u2212 x^2)dx + x(x^2y + 2x + y) = 0"

"y^2dx-x^2y^2dx+x^3ydy+2x^2dy+xydy=0\newline
ydx+xdy+x^3dy-x^2ydx+\frac{2x^2}{y}dy=0
\newline
d(xy)+x^4d(\frac{y}{x})+2x^2\frac{dy}{y}=0
\newline
(2Cxy+x^2+1)^2=-4Cx^2+x^4+2x^2+1, \text{ where C is some constant.}"